You have to find the measure of each angle..
The measure of one angle is twice the measure of its supplement.
How many angles are there? What is their relationship?
How can you use algebra, such as using the variable x, to help you.
The measure of one angle is 20 less than the measure of it's complement?

Answer by cleomenius(959)

x + 2x = 180
3x = 180
x = 60 degrees, one angle.
2x = 120 degrees, the second angle.
This does check.
There are two angles in the relationship.
We can use algebra by making one angle x and the second angle, the supplement, 2x.
=================================================================================
For the last part,
x + x + 20 = 90.
2x + 20 = 90
2x = 70
x = 35
x = 35 degrees, this is one angle.
35 + 20 = 55 degrees, this is the second angle.
35 degrees + 55 degrees = 90 degrees, this does check.
Cleomenius.